# SLAC-PUB-251124 January 2026

### v2

@@ 1 Introduction
These experimental advances ought to change our perspective on the study of particle physics. We now see clearly that the major outstanding questions about particle physics are questions about the interactions of the Higgs boson. We need to learn more about this particle, and we have the resources to gain that knowledge. We will learn much from the High Luminosity phase of the LHC, but, beyond this, we will be able to measure the properties of the Higgs boson with high precision from a future “Higgs Factory”, an $e^{+}e^{-}$ collilder operating at energies from the $Z$ boson resonance to the top quark threshold and above. This has been recognized in the latest strategy statements from all of the major world regions in particle physics. From the 2020 Update of the European Strategy for Particle Physics: “An electron-positron Higgs factor is the highest priority next collider.” [5]. From the 2022 report of the Japan Association of High Energy Physicists (JAHEP): “An $e^{+}e^{-}$ Higgs Factory is the most important next large-scale particle physics facility”[6]. From the 2023 US High Energy Physics Advisory Panel P5 report: “The next step is to use electron and positron beams to construct a Higgs factory, which would allow precision measurements of the Higgs boson properties and searches for exotic decays, possibly into dark matter” [8]. In China, the Circular Electron-Positron Collilder (CEPC) was named the highest priority next project in particle and nuclear physics in the first round of CAS deliberations [9], though eventually the CAS chose to endorse a less ambitious collider. Very recently, the European Strategy Group has recommended the FCC-ee, a circular Higgs factory, as the next flagship CERN project [10].
In this lecture, I review the importance of the measurements expected from Higgs factories and the levels of precision that they might achieve. Section 2 reviews the central role of the Higgs boson that motivates these measurements. In Sections 3–5, I discuss Higgs measurements in three different programs avaiable at Higgs factories. Section 3 discusses the program at the peak of the $e^{+}e^{-}\to HZ$ cross section at 240-250 GeV in the $e^{+}e^{-}$ center of mass. Section 4 discusses the program of precision measurements at the $Z$ resonance. Section 5 discusses measurements above the top quark threshold. All three sections touch on the question of the comparison of linear and circular $e^{+}e^{-}$ colliders. The European Strategy Group has now recommended the FCC-ee as the next CERN project, but still it is interesting to summarize what we have learned about this comparison in the past year of physics discussions. Section 6 discusses the fascinating technical opportunities that the Higgs factory physics program will provide. Section 7 will give some conclusions.
@@ 2 Importance of studying the Higgs boson with high precision
There is one more issue concerning the Higgs boson that needs a comment. This is the truly impressive job that the ATLAS and CMS experiments have done in measuring the couplings of the Higgs boson, using observed Higgs boson decays and also the production of $t\overline{t}H$ final states. The results as of the end of run 2 of the LHC are collected in Fig. 1, from CMS, and a similar figure from ATLAS, both published in Nature [3, 4]. This data demonstrates that the Higgs boson vacuum expectation value is the main source of mass for the particles shown. Many people interpret these beautiful results as proof that the observed Higgs boson precisely follows the Standard Model. To a great extent, this is true, but it is important to think more carefully.
@@ 3 Higgs measurements at threshold
Figure 3: Cross sections for Higgs boson production in $e^{+}e^{-}$ annhilation reactions as a function of CM energy, showing also the dominant Feynman diagrams for the reactions $e^{+}e^{-}\to ZH$ and $e^{+}e^{-}\to\nu\overline{\nu}H$.
@@
With this context, let’s begin our discussion of the expectations for the measurements of Higgs boson couplings – in particular, the couplings accessible in Higgs boson decay. The dominant cross sections for Higgs boson production at $e^{+}e^{-}$ colliders are Higgstrahlung, $e^{+}e^{-}\to ZH$, and $WW$ fusion, $e^{+}e^{-}\to\nu\overline{\nu}H$. The cross sections for these reactions in the CM energy range of Higgs factories is shown in Fig. 3. The cross sections for these two reactions cross over at about 470 GeV.
@@
It is straightforward to use this method to compare the expectations for the precision in Higgs coupling measurement between linear and circular colliders. In Table 1, I present the expected 1 $\sigma$ precisions (in %) for LCF, FCC-ee, and CEPC I assume the same detection efficiencies and background subtractions for all colliders, based on the full simulation studies done for ILC. This makes clearer the differences in the results that are due to the different run plans, that is, from the integrated luminosities at the various energy stages and from beam polarization for LCF. For LCF, I show the results for the initial stage at 250 GeV with 3 ab-1 of data and for the full program of including 8 ab-1 of data at 550 GeV. For FCC-ee and CEPC, the main sources of data are the runs at 240 GeV, but the estimates given include also anticipated running at the top quark threshold and just above. For FCC-ee, I use the run plan in [16], with 10.8 ab-1 of data at 240 GeV. For CEPC, I use the run plan in [19], with 2 detectors and a total of 20 ab-1 of data at 240 GeV. All estimates include expected results from HL-LHC on the $H\to\gamma\gamma$ and $H\to\mu^{+}\mu^{-}$ decays. The final results for the expected precisions are remarkably similar. In all cases, the uncertainties are dominated by statistical errors, with systematic effects well under control.
@@
A question often asked about Higgs coupling precision is: Yes, there are improvements over what is possible at LHC, but is it possible to access the effects of new particles? Given the success of the SM and the fact that, with a Higgs boson mass of 125 GeV, the SM could be valid up to the Planck scale, it is not possible to guarantee a discovery, either in direct particle searches or in precision studies. However, there are real oppportunites for discovery available. Some general classes of models giving sensitivity to new particles at the levels of precision given in Table 1 are reviewed in [28].
@@ 4 Study of the $Z$ boson at high statistics
On the second issue, I have much less confidence. The sources of uncertainty here are come from non-perturbative effects as they are modeled in the widely used Monte Carlo programs. This is not a matter of lacking statistics; what we lack is an accurate theoretical understanding of the non-perturbative stage of hadronization. It is noteworthy, and somewhat depressing, that we still estimate systematic errors from hadronization from the differences between the results of PYTHIA and HERWIG. Today, there is no strategy improve this situation. We should treat uncertainties from this source more conservatively.
@@
The “conservative” estimates of theory uncertainties are listed in Table 3, along with the statistical and experimental systematic uncertainties expected from the $Z$ pole programs of LCF and FCC-ee. It is important to note that these theory systematic uncertainties are large compared to the claimed experimental systematics, especially for the Tera-Z program. The resulting uncertainties on the $Z$ couplings to fermions are given in Table 4. In contrast, the the projections given in the Briefing Book for the uncertainties in Higgs couplings (Table 2), which are not at this very high level of precision, are quite robust to the choice of “conservative” or “aggressive” uncertainties. The differences are at the 20% level in the estimated uncertainty on the Higgs coupling for the $Z$ and $W$ and at the percent level for the other cases. This same robustness appears for the results on top quark interactions discussed in the following section.
@@
Even with the “conservative” theory systematic uncertainties, the power of the $Z$ pole measurements is very impressive. This is often presented as a table or graph of sensitivity scales $\Lambda$ for higher-dimension SMEFT operators. In (4), the coefficient of each dimension-6 operator is written as $c_{i}/\Lambda^{2}$, where $\Lambda^{2}$ is a mass scale. In an underlying BSM model, effects of new heavy particles generate corrections to the SM, and these are described by specific higher-dimension operators included in the SMEFT Lagrangian. In a fit to data with a Lagrangian including some number of these operators, the standard error $\sigma$ on each coefficient gives an idea of the masses of the new particles generating this operator to which that data set is sensitive. That scale would be roughtly $\Lambda\sim 1/\sqrt{\sigma}$ for operators generated by new strong interactions and $\Lambda/\sqrt{\alpha_{w}}\sim\sqrt{\sigma}$, given $\Lambda$ values a factor 6–10 smaller, for operators generated by electroweak radiative corrections.
@@ 5 Higgs measurements above the top quark threshold
Finally, I will discuss the Higgs factory program at energies above the top quark threshold. This region is often omitted from discusssions of the physics of Higgs factories, but in fact it allows additional important measurements whose results are essential to the Higgs story.
First, running at high energy will allow us to redo the measurements of the Higgs couplings listed in Table 1. As I pointed out above, the $WW$ fusion reaction becomes the dominant source of Higgs boson prodiuction above 400 GeV, so the new Higgs sample will be dominated by this new reaction, with different experimental issues and sources of uncertainty. This will be an important check on any anomalies discovered in the program at the Higgs threshold. Assuming compatibility of the two sets of results, the improvement in the uncertainties on Higgs couplings has already been shown in Table 1.
@@
This Table requires some explanation. The operators included in the left-hand column are: the 3 operators that shift the left- and right-handed top quark electroweak couplings, the operator that shifts the top quark Yukawa coupling, the 5 operators that provide 4-fermion $e^{+}e^{-}t\overline{t}$ contact interactions, and the 2 operators that shift the top quark $W$ and $Z$ magnetic moments. All of these operators can be induced by BSM top quark physics. I have discussed the shift in the electroweak couplings above. The 4-fermion operators can be induced by constituent exchange in composite top quark models, or by the Kaluza-Klein vector bosons in a higher-dimensional model of Higgs compositeness. Four-fermion $e^{+}e^{-}t\overline{t}$ operators can also be induced as NLO corrections to order-1 4-top contact interactions. It is important to recognize that the SM is violated by BSM effects, more important to demonstrate that this effect is specifically associated with the top quark, and still more valuable to distinguish between the electroweak couplings and 4-fermion effects, which are characteristic of different classes of models.
@@
The single-operator sensitivities in the first two columns are already very impressive. These are mainly obtained from the $Z$ pole programs of the LCF and FCC-ee. As I noted in Section 4, the Tera-Z program gives higher sensitivities, but only to the extent that the uncertainties presented Table 4 are smaller. The next two columns include the first stages beyond the top quark threshold. The sensitivites from direct $t\overline{t}$ observations surpass those from the $Z$ pole measurements, and more substantially as the CM energy is raised. The specificity of the operator identifications is still low, because, with only one CM energy, there is a degeneracy in the effects from the top quark electroweak modifications and the top 4-fermion couplings. However, the electroweak modifications to the cross section stay at a constant size as a function of CM energy, while the effects of 4-fermion operators increase as $E_{CM}^{2}$. Then, adding a data set at 1 TeV clearly distinguishes the two effects, giving marked increases in the ability of the fit to pick out the specific top quark operator responsible. Also, to the extent that the new physics induces 4-fermion interactions, the 1 TeV data substantially improves the overall sensitivity to BSM physics.
At the CM energies where top quark observations dominate the SMEFT fit results, those results again become relatively insensitive to theory systematics, giving very similar results from the “conservative” and “aggressive” assumptions discussed in Section 4. I commented there that observations of tiny deviations in the precision electroweak observables may not have sufficient power to convince skeptics that the SM is in fact violated. For BSM physics associated with the third generation, observation of deviations from the SM predictions of the top quark interactions would strongly confirm any effect suggested by the precision electroweak program. This statement generalizes to other source of BSM physics, since high energy $e^{+}e^{-}$ experiments will look for deviations in the complete set of final state available through 4-fermion contact interactions.
@@ 6 The challenge for now
The construction of a Higgs factory to begin operation just after the completion of the HL-LHC program is important for our understanding of physics, but it is even more important for the early career scientists in our community. The HL-LHC, planned to operate in 2030, offers exciting opportunites for the discovery of new particles, especially for color-neutral states such as extended Higgs states and electroweakinos with masses in the range of 1 TeV. But we can already see the limits of this program and the need to look beyond it. A new major collider takes 10 years to construct. This will be the dominant machine at the peak of your career. It is time to start thinking about it now.
This is going to be fun. The environment of an $e^{+}e^{-}$ collider is very different from that of the LHC. There is minimal radiation load on the detectors, no underlying event, no pileup. The $e^{+}e^{-}$ annhilation events are simpler than those at the LHC. Not only are the particle multiplicities lower, but there is no forward fragmentation lost down the beam pipe. This is an environment in which it is possible to make vertex, tracking, and calorimetry measurements that are close to perfect. But – this point is very important – this does not make experimentation easier, because great improvements in event reconstruction are need to satisfy the goals of high precision measurements.
@@
An alternative technology would be dual readout calorimetry, in which one would collect both scintillation and Cherenkov light signals from calorimetry modules, allowing separation of the charged and neutral components of a hadronic shower [46]. This method is now undergoing rapid development, with the inclusion of longitudinal segmentation, timing, and particle flow analysis strategies [47, 48]. Up to now, incorporation of Machine Learning in either technique has not led to large improvements in resolution, but it is surely just a matter of time before those ideas have their effect.
@@
This brings up a more general question that I feel is very important. So far, all of the AI applications to event analysis have been done in the context of existing designs and hardware, both for LHC and for Higgs factories. But to develop a develop a future detector, as AI gives us more capabilities, the hardware design should co-evolve with the analysis software to use the new methods most effectively. Detectors for the 2040’s will have some level of intelligence on every sensor. Does this push us to higher granularity to gain more information or to lower granularity, with local analysis of the waveforms produced by the sensors? How will we read out highly intelligent sensors? How will we incorporate timing information? How to these solutions interact with more traditional restrictions to minimize material and heating? The answers to these questions may well lead us to radically new ideas about detector design and performance.

### v3

@@ 1 Introduction
These experimental advances ought to change our perspective on the study of particle physics. We now see clearly that the major outstanding questions about particle physics are questions about the interactions of the Higgs boson. We need to learn more about this particle, and we have the resources to gain that knowledge. We will learn much from the High Luminosity phase of the LHC, but, beyond this, we will be able to measure the properties of the Higgs boson with high precision from a future “Higgs Factory”, an $e^{+}e^{-}$ collider operating at energies from the $Z$ boson resonance to the top quark threshold and above. This has been recognized in the latest strategy statements from all of the major world regions in particle physics. From the 2020 Update of the European Strategy for Particle Physics: “An electron-positron Higgs factor is the highest priority next collider.” [5]. From the 2022 report of the Japan Association of High Energy Physicists (JAHEP): “An $e^{+}e^{-}$ Higgs Factory is the most important next large-scale particle physics facility”[6]. From the 2023 US High Energy Physics Advisory Panel P5 report: “The next step is to use electron and positron beams to construct a Higgs factory, which would allow precision measurements of the Higgs boson properties and searches for exotic decays, possibly into dark matter” [8]. In China, the Circular Electron-Positron Collider (CEPC) was named the highest priority next project in particle and nuclear physics in the first round of CAS deliberations [9], though eventually the CAS chose to endorse a less ambitious collider. Very recently, the European Strategy Group has recommended the FCC-ee, a circular Higgs factory, as the next flagship CERN project [10].
In this lecture, I review the importance of the measurements expected from Higgs factories and the levels of precision that they might achieve. Section 2 reviews the central role of the Higgs boson that motivates these measurements. In Sections 3–5, I discuss Higgs measurements in three different programs available at Higgs factories. Section 3 discusses the program at the peak of the $e^{+}e^{-}\to HZ$ cross section at 240-250 GeV in the $e^{+}e^{-}$ center of mass. Section 4 discusses the program of precision measurements at the $Z$ resonance. Section 5 discusses measurements above the top quark threshold. All three sections touch on the question of the comparison of linear and circular $e^{+}e^{-}$ colliders. The European Strategy Group has now recommended the FCC-ee as the next CERN project, but still it is interesting to summarize what we have learned about this comparison in the past year of physics discussions. Section 6 discusses the fascinating technical opportunities that the Higgs factory physics program will provide. Section 7 will give some conclusions.
@@ 2 Importance of studying the Higgs boson with high precision
There is one more issue concerning the Higgs boson that needs a comment. This is the truly impressive job that the ATLAS and CMS experiments have done in measuring the couplings of the Higgs boson, using observed Higgs boson decays and also the production of $t\overline{t}H$ final states. The results as of the end of run 2 of the LHC are collected in Fig. 1, from CMS, and a similar figure from ATLAS, both published in Nature [3, 4]. These data demonstrate that the Higgs boson vacuum expectation value is the main source of mass for the particles shown. Many people interpret these beautiful results as proof that the observed Higgs boson precisely follows the Standard Model. To a great extent, this is true, but it is important to think more carefully.
@@ 3 Higgs measurements at threshold
Figure 3: Cross sections for Higgs boson production in $e^{+}e^{-}$ annihilation reactions as a function of CM energy, showing also the dominant Feynman diagrams for the reactions $e^{+}e^{-}\to ZH$ and $e^{+}e^{-}\to\nu\overline{\nu}H$.
@@
With this context, let’s begin our discussion of the expectations for the measurements of Higgs boson couplings – in particular, the couplings accessible in Higgs boson decay. The dominant cross sections for Higgs boson production at $e^{+}e^{-}$ colliders are Higgstrahlung, $e^{+}e^{-}\to ZH$, and $WW$ fusion, $e^{+}e^{-}\to\nu\overline{\nu}H$. The cross sections for these reactions in the CM energy range of Higgs factories are shown in Fig. 3. The cross sections for these two reactions cross over at about 470 GeV.
@@
It is straightforward to use this method to compare the expectations for the precision in Higgs coupling measurement between linear and circular colliders. In Table 1, I present the expected 1 $\sigma$ precisions (in %) for LCF, FCC-ee, and CEPC. I assume the same detection efficiencies and background subtractions for all colliders, based on the full simulation studies done for ILC. This makes clearer that the differences in the results that are due to the different run plans, that is, from the integrated luminosities at the various energy stages and from beam polarization for LCF. For LCF, I show the results for the initial stage at 250 GeV with 3 ab-1 of data and for the full program of including 8 ab-1 of data at 550 GeV. For FCC-ee and CEPC, the main sources of data are the runs at 240 GeV, but the estimates given include also anticipated running at the top quark threshold and just above. For FCC-ee, I use the run plan in [16], with 10.8 ab-1 of data at 240 GeV. For CEPC, I use the run plan in [19], with 2 detectors and a total of 20 ab-1 of data at 240 GeV. All estimates include expected results from HL-LHC on the $H\to\gamma\gamma$ and $H\to\mu^{+}\mu^{-}$ decays. The final results for the expected precisions are remarkably similar. In all cases, the uncertainties are dominated by statistical errors, with systematic effects well under control.
@@
A question often asked about Higgs coupling precision is: Yes, there are improvements over what is possible at LHC, but is it possible to access the effects of new particles? Given the success of the SM and the fact that, with a Higgs boson mass of 125 GeV, the SM could be valid up to the Planck scale, it is not possible to guarantee a discovery, either in direct particle searches or in precision studies. However, there are real opportunities for discovery available. Some general classes of models giving sensitivity to new particles at the levels of precision given in Table 1 are reviewed in [28].
@@ 4 Study of the $Z$ boson at high statistics
On the second issue, I have much less confidence. The sources of uncertainty here come from non-perturbative effects as they are modeled in the widely used Monte Carlo programs. This is not a matter of lacking statistics; what we lack is an accurate theoretical understanding of the non-perturbative stage of hadronization. It is noteworthy, and somewhat depressing, that we still estimate systematic errors from hadronization from the differences between the results of PYTHIA and HERWIG. Today, there is no strategy improve this situation. We should treat uncertainties from this source more conservatively.
@@
The “conservative” estimates of theory uncertainties are listed in Table 3, along with the statistical and experimental systematic uncertainties expected from the $Z$ pole programs of LCF and FCC-ee. It is important to note that these theory systematic uncertainties are large compared to the claimed experimental systematics, especially for the Tera-Z program. The resulting uncertainties on the $Z$ couplings to fermions are given in Table 4. In contrast, the projections given in the Briefing Book for the uncertainties in Higgs couplings (Table 2), which are not at this very high level of precision, are quite robust to the choice of “conservative” or “aggressive” uncertainties. The differences are at the 20% level in the estimated uncertainty on the Higgs coupling for the $Z$ and $W$ and at the percent level for the other cases. This same robustness appears for the results on top quark interactions discussed in the following section.
@@
Even with the “conservative” theory systematic uncertainties, the power of the $Z$ pole measurements is very impressive. This is often presented as a table or graph of sensitivity scales $\Lambda$ for higher-dimension SMEFT operators. In (4), the coefficient of each dimension-6 operator is written as $c_{i}/\Lambda^{2}$, where $\Lambda^{2}$ is a mass scale. In an underlying BSM model, effects of new heavy particles generate corrections to the SM, and these are described by specific higher-dimension operators included in the SMEFT Lagrangian. In a fit to data with a Lagrangian including some number of these operators, the standard error $\sigma$ on each coefficient gives an idea of the masses of the new particles generating this operator to which that data set is sensitive. That scale would be roughly $\Lambda\sim 1/\sqrt{\sigma}$ for operators generated by new strong interactions and $\Lambda/\sqrt{\alpha_{w}}\sim\sqrt{\sigma}$, for operators generated by electroweak radiative corrections, giving $\Lambda$ values a factor 6–10 smaller.
@@ 5 Higgs measurements above the top quark threshold
Finally, I will discuss the Higgs factory program at energies above the top quark threshold. This region is often omitted from discussions of the physics of Higgs factories, but in fact it allows additional important measurements whose results are essential to the Higgs story.
First, running at high energy will allow us to redo the measurements of the Higgs couplings listed in Table 1. As I pointed out above, the $WW$ fusion reaction becomes the dominant source of Higgs boson production above 400 GeV, so the new Higgs sample will be dominated by this new reaction, with different experimental issues and sources of uncertainty. This will be an important check on any anomalies discovered in the program at the Higgs threshold. Assuming compatibility of the two sets of results, the improvement in the uncertainties on Higgs couplings has already been shown in Table 1.
@@
This Table requires some explanation. The operators included in the left-hand column are: the 3 operators that shift the left- and right-handed top quark electroweak couplings, the operator that shifts the top quark Yukawa coupling, the 5 operators that provide 4-fermion $e^{+}e^{-}t\overline{t}$ contact interactions, and the 2 operators that shift the top quark $W$ and $Z$ magnetic moments. All of these operators can be induced by BSM top quark physics. I have discussed the shift in the electroweak couplings above. The 4-fermion operators can be induced by constituent exchange in composite top quark models, or by the Kaluza-Klein vector bosons in a higher-dimensional model of Higgs compositeness. Four-fermion $e^{+}e^{-}t\overline{t}$ operators can also be induced as NLO corrections to order-1 4-top contact interactions.
@@
The single-operator sensitivities in the first two columns are already very impressive. These are mainly obtained from the $Z$ pole programs of the LCF and FCC-ee. As I noted in Section 4, the Tera-Z program gives higher sensitivities, but only to the extent that the uncertainties presented Table 4 are smaller. The next two columns include the first stages beyond the top quark threshold. The sensitivities from direct $t\overline{t}$ observations surpass those from the $Z$ pole measurements, and more substantially as the CM energy is raised. The specificity of the operator identifications is still low, because, with only one CM energy, there is a degeneracy in the effects from the top quark electroweak modifications and the top 4-fermion couplings. However, the electroweak modifications to the cross section stay at a constant size as a function of CM energy, while the effects of 4-fermion operators increase as $E_{CM}^{2}$. Then, adding a data set at 1 TeV clearly distinguishes the two effects, giving marked increases in the ability of the fit to pick out the specific top quark operator responsible. Also, to the extent that the new physics induces 4-fermion interactions, the 1 TeV data substantially improves the overall sensitivity to BSM physics.
This study addresses a number of goals. It is important to recognize that the SM is violated by BSM effects, more important to demonstrate that this effect is specifically associated with the top quark, and still more valuable to distinguish between the electroweak couplings and 4-fermion effects, which are characteristic of different classes of models. Table 5 shows that the program of high-energy running of an $e^{+}e^{-}$ collider will address all of these questions.
At the CM energies where top quark observations dominate the SMEFT fit results, those results again become relatively insensitive to theory systematics, giving very similar results from the “conservative” and “aggressive” assumptions discussed in Section 4. I commented there that observations of tiny deviations in the precision electroweak observables may not have sufficient power to convince skeptics that the SM is in fact violated. For BSM physics associated with the third generation, observation of deviations from the SM predictions of the top quark interactions would strongly confirm any effect suggested by the precision electroweak program. This statement generalizes to other source of BSM physics, since high energy $e^{+}e^{-}$ experiments will look for deviations in the complete set of final states available through 4-fermion contact interactions.
@@ 6 The challenge for now
The construction of a Higgs factory to begin operation just after the completion of the HL-LHC program is important for our understanding of physics, but it is even more important for the early career scientists in our community. The HL-LHC, planned to operate in 2030, offers exciting opportunities for the discovery of new particles, especially for color-neutral states such as extended Higgs states and electroweakinos with masses in the range of 1 TeV. But we can already see the limits of this program and the need to look beyond it. A new major collider takes 10 years to construct. This will be the dominant machine at the peak of your career. It is time to start thinking about it now.
This is going to be fun. The environment of an $e^{+}e^{-}$ collider is very different from that of the LHC. There is minimal radiation load on the detectors, no underlying event, no pileup. The $e^{+}e^{-}$ annihilation events are simpler than those at the LHC. Not only are the particle multiplicities lower, but there is no forward fragmentation lost down the beam pipe. This is an environment in which it is possible to make vertex, tracking, and calorimetry measurements that are close to perfect. But – this point is very important – this does not make experimentation easier, because great improvements in event reconstruction are need to satisfy the goals of high precision measurements.
@@
An alternative technology would be dual readout calorimetry, in which one would collect both scintillation and Cherenkov light signals from calorimetry modules, allowing separation of the electromagnetic and non-electromagnetic components of a hadronic shower [46]. This method is now undergoing rapid development, with the inclusion of longitudinal segmentation, timing, and particle flow analysis strategies [47, 48]. Up to now, incorporation of Machine Learning in either technique has not led to large improvements in resolution, but it is surely just a matter of time before those ideas have their effect.
@@
This brings up a more general question that I feel is very important. So far, all of the AI applications to event analysis have been done in the context of existing designs and hardware, both for LHC and for Higgs factories. But to develop a develop a future detector, as AI gives us more capabilities, the hardware design should co-evolve with the analysis software to use the new methods most effectively. Detectors for the 2040’s will have some level of intelligence on every sensor. Does this push us to higher granularity to gain more information or to lower granularity, with local analysis of the waveforms produced by the sensors? How will we read out highly intelligent sensors? How will we incorporate timing information? How do these solutions interact with more traditional restrictions to minimize material and heating? The answers to these questions may well lead us to radically new ideas about detector design and performance.